OpenStudy (anonymous):

is y(t)=sint + 4cos2πt periodic?

OpenStudy (tkhunny):

What is the definition of "Periodic"?

OpenStudy (tkhunny):

Is that \(\cos^{2}(\pi t)\;or\;\cos(2\pi t)\)

OpenStudy (tkhunny):

Is y(t) = y(t+k) for ANY k?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."